OpenStudy (anonymous):

Find the shortest distance from the line 2x+6y=3 and passing through the point (9,-7)

OpenStudy (mertsj):

That would be the perpendicular distance so you must first find the equation of the perpendicular through( 9,-7) that has slope 3.

OpenStudy (mertsj):

Do that and we'll go from there.

OpenStudy (anonymous):

can yu show me how to do it?

OpenStudy (anonymous):

If memory serves, \(\huge \frac{ 2(9)+6(-7)-3 } { \pm \sqrt{2^2+6^2 } }\)

OpenStudy (mertsj):

Can you find the equation of a line whose slope is 3 and contains the point (9,-7)?

OpenStudy (mertsj):

Oh. I see that Fool wants to help you now so I'll turn you over to him.

OpenStudy (anonymous):

lol yu can still help me

OpenStudy (anonymous):

im not sure if yur suppose to use the quadratic formula for this

OpenStudy (mimi_x3):

THe formula for this, i think: \[\large \frac{ax_{1}+by_{1}+c}{\sqrt{a^{2}+b^{2}}} \]

OpenStudy (anonymous):

there is a formula for it

OpenStudy (anonymous):

The proof requires representing the lines in normal form and then finding the perpendicular distances.

OpenStudy (anonymous):

Mimi, your formula is close but not apt, it should be Either, \[ \large \pm\frac{ax_{1}+by_{1}+c}{\sqrt{a^{2}+b^{2}}} \] or \[ \large \frac{|ax_{1}+by_{1}+c|}{\sqrt{a^{2}+b^{2}}} \] Recall, Euclidean distance cannot be negative :)

OpenStudy (mimi_x3):

lol, yeah i forgot my bad.

OpenStudy (anonymous):

no worries, as long as you give me medal :P

OpenStudy (mimi_x3):

obssessed of medals huh :P

OpenStudy (anonymous):

whats the point of medals? do yu get nythin

OpenStudy (anonymous):

hehe, not really it's fun though.

OpenStudy (mimi_x3):

You get nothing, just a level, it just a waste of time anyway :P

OpenStudy (anonymous):

Why should I answer you? Will I get anything? What is the purpose of anything in life? Life is as meaningless itself as this discussion :P

OpenStudy (anonymous):

LMAO ^

OpenStudy (mimi_x3):

Lol, life is cruel.

OpenStudy (anonymous):

;)